### v2

# How Conflict Aversion Can Enable Authoritarianism: An Evolutionary Dynamics Approach

Abstract
We use evolutionary game theory to examine how conflict-averse centrism can facilitate authoritarian success in polarized political conflicts. Such conflicts are often asymmetric: authoritarian actors can employ norm-breaking or coercive tactics, while democratic resistance faces stronger normative constraints on acceptable behavior. Yet formal models typically treat sides symmetrically and rarely examine conflict-averse behavior. Drawing on empirical research on protest backlash, civility norms, and authoritarian resilience, we model these dynamics as a three-strategy evolutionary game. This framework yields two outcomes—cyclic authoritarian resurgence through a heteroclinic cycle and a stable centrist–authoritarian coalition excluding resistance—depending on confrontation responses. We demonstrate how an established dynamical framework with empirically grounded behavioral assumptions clarifies conditions under which conflict aversion can diminish the effectiveness of democratic resistance.
1 Introduction
The contemporary political landscape in the United States is often described as “polarized,” marked by intense partisan animus and divergent political worldviews (HetheringtonWeiler2009; Iyengar2012; Mason2018). This conventional framing treats opposing poles as normatively and strategically symmetric. Yet many real-world conflicts involve asymmetric aims: movements seeking democratic or egalitarian expansion face actors pursuing exclusionary or authoritarian objectives. Research on democratic backsliding, asymmetric party development, and social-movement repression shows that these poles differ systematically in goals, tactics, and institutional incentives (LevitskyZiblatt2018; Bermeo2016; PiersonSchickler2020; Earl2011).
Empirical work highlights a third behavioral category that interacts with, but does not map neatly onto, either pole: individuals driven more by conflict aversion than by policy stakes. This aversion manifests in several well-documented ways. Experiments and observational studies show that substantial segments of the public penalize tactics they see as extreme, disruptive, or otherwise norm-violating, even when they sympathize with the underlying goals (FeinbergWiller2020; Wasow2020). Many self-identified independents are motivated as much by a desire to avoid partisan conflict as by ideological moderation (KlarKrupnikov2016). Classic work on political discussion indicates that conflict-averse individuals often withdraw from settings where disagreement is expected, reducing political engagement even when they care about the issues (e.g., Mutz2006).
People who dislike confrontation may selectively exit contexts explicitly labeled as “political,” yet remain willing to discuss the same issues when conflict is downplayed (GroenendykKrupnikovRyanConnors2025). These individuals are not ideologically centrist so much as behaviorally conflict-averse. Their responses to confrontation differ systematically from both resistance and authoritarian strategies, making them a distinct type rather than a midpoint on a left–right spectrum.
Although such actors often remain muted or invisible in media discourse (Bail2021), they constitute a large share of the mass public (Mutz2018) and shape the strategic landscape in which opposing forces interact. Despite this empirical importance, conflict aversion rarely appears as a formalized behavioral strategy in theoretical models. Most models of polarization treat individuals as points in a continuous metric space and define polarization as a distributional property of those points (FriedkinJohnsen1999; Deffuant2000; HegselmannKrause2002; FlacheMacy2011; TurnerSmaldino2018; Duggins2017; FlacheEtAl2017). These frameworks model contending poles symmetrically with respect to normative content and strategic objectives: their influence equations do not distinguish differences in normative content or strategic aims. Related evolutionary models that include centrist or moderate strategies likewise interpret centrism as an ideological midpoint rather than a preference for minimizing confrontation (e.g., ArceSandler2003; ShortEtAl2017).
These patterns suggest that conflict aversion functions as a distinct strategic type with its own payoff structure, motivating a game-theoretic approach that treats centrists’ responses to resistance and authoritarianism asymmetrically. We therefore develop a minimal evolutionary-game model with three behavioral strategies: authoritarianism or fascism, resistance to authoritarian encroachment, and conflict-averse centrism that imposes reputational or social penalties on visible confrontation. We encode strategy payoffs in a three-by-three game and study their population dynamics using replicator dynamics (TaylorJonker1978; HofbauerSigmund1998).
We then analyze two empirically motivated strategic regimes: a cyclic resurgence regime in which resistance defeats authoritarianism in isolation but conflict-averse centrism generates a heteroclinic cycle, and a coalition regime in which mutually reinforcing centrists and authoritarians stabilize against resistance. Across these regimes, empirically grounded assumptions about resistance, conflict aversion, and authoritarian norm-breaking overturn the expectation that pairwise dominance of resistance over authoritarianism guarantees long-run success.
Mathematically, our analysis draws on classical results for three-strategy replicator systems, including the characterization of interior and boundary equilibria, conditions for cyclic dominance, and the structure of heteroclinic cycles (Zeeman1980; Bomze1983; HofbauerSigmund1998). These results clarify the dynamical possibilities that arise when conflict aversion is modeled as a strategic type. Our contribution is to show that empirically motivated payoffs naturally generate two distinct mechanisms by which authoritarianism can prevail. Taken together, these mechanisms illuminate how conflict-averse centrism can unintentionally facilitate authoritarian success.
2 Background
To situate our contribution, we review relevant work on opinion dynamics and evolutionary games. Formal models of opinion formation typically represent individuals as adjusting their expressed views in response to interaction partners. Linear-averaging frameworks show how stable disagreement can persist under repeated mutual influence (FriedkinJohnsen1999). Bounded-confidence models restrict interaction to partners whose opinions lie within a specified tolerance, producing consensus, clustering, or fragmentation depending on parameters and initial conditions (Deffuant2000; HegselmannKrause2002). Extensions incorporate network topology, repulsive influence, extremist effects, and—in some cases—stochastic elements to explain durable polarization and the emergence of opinion subcultures (FlacheMacy2011; TurnerSmaldino2018; Duggins2017; FlacheEtAl2017).
These frameworks share key structural features: opinions are represented as points in a metric space, and polarization is a distributional property of those points. Interacting poles are modeled symmetrically in the sense that the equations governing influence do not distinguish normative content or strategic objectives. Even when heterogeneity in responsiveness is included, the models do not treat behavior shaped primarily by conflict aversion as a distinct type. Thus, they illuminate how clustering emerges from local influence rules but offer limited insight into how conflict-averse behavior might alter asymmetric strategic contests. Evolutionary models with multiple ideological or attitudinal positions likewise treat “moderate” types as ideologically intermediate rather than as actors motivated by minimizing confrontation (e.g., ArceSandler2003; ShortEtAl2017).
A parallel empirical literature documents the prevalence and behavioral consequences of conflict aversion: individuals penalize confrontational tactics (FeinbergWiller2020; Wasow2020), avoid partisan conflict (KlarKrupnikov2016), withdraw from expected disagreement (e.g., Mutz2006), and selectively exit explicitly political contexts while engaging when conflict cues are muted (GroenendykKrupnikovRyanConnors2025). These findings indicate that conflict-averse centrism is both coherent and influential, yet it has no formal analog in standard models of polarization.
Evolutionary game theory provides a complementary framework in which strategic performance depends on payoffs that vary with population composition and interaction patterns. Empirical research on social influence shows that behaviors can spread through social interaction and network effects (e.g., Rogers2003; Centola2010; Sinclair2012). This makes evolutionary games particularly well suited for modeling how confrontational or conciliatory strategies spread at the population level.
Three-strategy games can exhibit cyclic dominance: $A$ outperforms $B$, $B$ outperforms $C$, and $C$ outperforms $A$, where $A$, $B$, and $C$ denote behavioral strategies whose population shares evolve over time. For broader overviews of cyclic-dominance phenomena in evolutionary game and ecological models, see Szolnoki2014. In continuous-time models related to the replicator equation, such as the three-species Lotka–Volterra system studied by MayLeonard1975, cyclic dominance can generate interior limit cycles. For the standard three-strategy replicator dynamics, however, the global phase portraits are more constrained: classical analyses show that no isolated asymptotically stable periodic orbits occur in the simplex interior and that non-equilibrium recurrent behavior, when it arises, does so through heteroclinic cycles on the boundary for appropriate sign patterns (Zeeman1980; Bomze1983; HofbauerSigmund1998). These structural results emphasize that pairwise dominance relations do not uniquely determine global outcomes, and that a third strategy can qualitatively reorganize long-run dynamics. This motivates examining whether conflict-averse centrism plays a mediating role when interacting with resistance and authoritarian strategies, and whether its behavioral incentives can generate indirect or unintuitive pathways through which authoritarian strategies prevail.
3 Modeling
We consider a population composed of three behavioral strategies: resistance ($R$), authoritarian fascism ($F$, “fascism” for short), and conflict-averse centrism ($C$). Let $x(t)=(x_{R},x_{C},x_{F})$ denote the population frequencies of these strategies at time $t$, constrained to the simplex
$\Delta^{2}=\{x\in\mathbb{R}^{3}_{\geq 0}:x_{R}+x_{C}+x_{F}=1\}.$ (1)
An edge of the simplex refers to the one-dimensional boundary on which exactly two strategies are present.
A payoff matrix $A\in\mathbb{R}^{3\times 3}$ encodes strategic interactions: $A_{ij}$ is the payoff to a focal individual using (row) strategy $i$ when paired with an opponent using (column) strategy $j$. In population state $x$, the expected payoff to strategy $i$ is
$\pi_{i}(x)=(Ax)_{i}=\sum_{j\in\{R,C,F\}}A_{ij}x_{j},$ (2)
and the population-average payoff is
$\bar{\pi}(x)=x^{\top}Ax.$ (3)
Frequencies evolve according to replicator dynamics (TaylorJonker1978; HofbauerSigmund1998),
$\dot{x}_{i}=x_{i}\big[(Ax)_{i}-x^{\top}Ax\big],\qquad i\in\{R,C,F\},$ (4)
so a strategy increases in frequency when it earns above-average payoff.
The replicator dynamics possess a useful invariance that allows us to simplify $A$ without loss of generality. If we add a constant $d_{j}$ to every entry in column jj—that is, to all payoffs against opponent strategy jj—the modified matrix $A^{\prime}$ satisfies
$(A^{\prime}x)_{i}=(Ax)_{i}+\sum_{j}d_{j}x_{j},$ (5)
which shifts all strategy payoffs by the same amount $\sum_{j}d_{j}x_{j}$. Because the dynamics depend only on payoff differences, $\pi_{i}(x)-\bar{\pi}(x)$ is unchanged. Thus column shifts leave the dynamics invariant, and we may set the diagonal entries to zero without loss of generality by choosing $d_{j}=-A_{jj}$.
This invariance also preserves invasion fitness. Let $e_{j}$ denote the vertex at which all individuals play strategy $j$. A rare mutant strategy $i$ is said to invade strategy $j$ if its initial per-capita growth rate is positive when introduced at infinitesimal frequency into a resident population at $e_{j}$; this is a local statement about the stability of the vertex. The invasion fitness of $i$ against $j$ is
$\frac{\dot{x}_{i}}{x_{i}}\Big|_{x=e_{j}}=\pi_{i}(e_{j})-\pi_{j}(e_{j})=A_{ij}-A_{jj},$ (6)
which is unaffected by column shifts. Under diagonal normalization this becomes simply $A_{ij}$, so the sign of $A_{ij}$ alone determines whether $i$ can invade $j$. Off-diagonal entries therefore directly encode invasion advantages or disadvantages.
Adopting the diagonal normalization $A_{ii}=0$, the replicator dynamics can be written explicitly as
$\dot{x}_{R}$ $=x_{R}\,(A_{RC}x_{C}+A_{RF}x_{F}-\bar{\pi}),$ (7)
$\dot{x}_{C}$ $=x_{C}\,(A_{CR}x_{R}+A_{CF}x_{F}-\bar{\pi}),$ (8)
$\dot{x}_{F}$ $=x_{F}\,(A_{FR}x_{R}+A_{FC}x_{C}-\bar{\pi}),$ (9)
with
$\bar{\pi}=x_{R}(A_{RC}x_{C}+A_{RF}x_{F})+x_{C}(A_{CR}x_{R}+A_{CF}x_{F})+x_{F}(A_{FR}x_{R}+A_{FC}x_{C}).$ (10)
To connect the model to observable political behavior, we motivate the signs of the key payoff entries using findings from political science and political psychology. Centrism in this model prioritizes conflict reduction and civility over direct confrontation, consistent with evidence that many citizens penalize behavior perceived as uncivil, disruptive, or polarizing (Mutz2006; KlarKrupnikov2016; FeinbergWiller2020; GroenendykKrupnikovRyanConnors2025). We therefore assume that when $R$ and $C$ interact, resistance incurs a reputational or social cost while centrism gains a signaling benefit:
$A_{RC}=-a_{R},\qquad A_{CR}=b_{C},\qquad a_{R},b_{C}>0.$ (11)
This sign pattern is held fixed throughout.
Interactions between centrism and authoritarian fascism vary across environments. In some settings, centrists incur costs when confronting authoritarian norm-breaking, while authoritarians benefit from exploiting centrists’ commitments to civility and restraint (LevitskyZiblatt2018; BoykoffBoykoff2004; Bermeo2016; Gerschewski2021; LandauDixon2020). In others, centrists may benefit from accommodating authoritarian actors, for example through perceived stability, access, or protection of the status quo (e.g., Paxton2004; Berman2019; JostFedericoNapier2009). To describe strategic relations along a two-strategy edge, we write $i\succ j$ when strategy $i$ can invade strategy $j$ but not vice versa, and $i\leftrightarrow j$ when both strategies can invade each other. Empirical accounts of authoritarian resilience and centrist accommodation indicate two relevant configurations for the CCFF edge: fascist dominance ($F\succ C$) and mutual invadability ($C\leftrightarrow F$). The remaining logical case, $C\succ F$, is not the focus here.
For the RRFF interaction, resistance gains a payoff advantage over fascism ($A_{RF}=b_{R}>0$) while fascism incurs a penalty ($A_{FR}=-a_{F}<0$), consistent with $R\succ F$. For the CCFF interaction, the two regimes we study assume either that fascism exploits centrists ($A_{CF}=-a_{C}<0$ and $A_{FC}=b_{F}>0$, yielding $F\succ C$) or that $C$ and $F$ benefit simultaneously ($A_{CF}=c_{C}>0$ and $A_{FC}=c_{F}>0$, yielding $C\leftrightarrow F$).
The central motivating question of this paper is how fascism can prevail even when resistance is intrinsically stronger in direct confrontation. We therefore restrict attention to environments in which resistance defeats fascism in isolation ($R\succ F$), a pattern supported by research on the effectiveness of nonviolent resistance (ChenowethStephan2011; Wasow2020). Within this empirically grounded baseline, the two configurations along the CCFF edge just described give rise to two distinct strategic environments. In the next section, we analyze the dynamics induced by these environments and show how conflict-averse centrism can enable authoritarianism to dominate long-run outcomes despite resistance’s inherent advantage.
4 Results
We now determine which global dynamics arise from the two empirically grounded sign configurations analyzed below. For three-strategy replicator dynamics, trajectories may converge to vertex equilibria, edge equilibria, a unique interior equilibrium, or a boundary heteroclinic cycle, and isolated periodic orbits cannot be stable (Zeeman1980; Bomze1983; HofbauerSigmund1998). Our aim is not to assume any attractor in advance but to determine which of these outcomes follow from the payoff sign patterns implied by our behavioral assumptions. As we will show, two distinct global behaviors emerge directly from those signs: one in which the induced dominance cycle produces a heteroclinic orbit on which trajectories spend asymptotically disproportionate time near the fascist vertex, and one in which mutual invadability along the CCFF edge yields a stable centrist–fascist coalition that excludes resistance.
We begin with the sign configuration in which resistance defeats fascism in isolation, centrism penalizes resistance, and fascism exploits centrism. These behavioral assumptions are encoded in the diagonal-normalized payoff matrix
$A^{(1)}=\begin{pmatrix}0&-a_{R}&b_{R}\\ b_{C}&0&-a_{C}\\ -a_{F}&b_{F}&0\end{pmatrix}$, (12)
where $a_{R},a_{C},a_{F},b_{R},b_{C},b_{F}>0$. Expected payoffs follow from matrix multiplication:
$\pi_{R}(x)=-a_{R}x_{C}+b_{R}x_{F},\qquad\pi_{C}(x)=b_{C}x_{R}-a_{C}x_{F},\qquad\pi_{F}(x)=-a_{F}x_{R}+b_{F}x_{C}.$ (13)
On the RRFF edge, $x_{C}=0$ and $x_{F}=1-x_{R}$. Substituting into (13) yields
$\pi_{R}=b_{R}(1-x_{R}),\qquad\pi_{F}=-a_{F}x_{R},$ (14)
and therefore
$\pi_{R}-\pi_{F}=b_{R}+x_{R}(a_{F}-b_{R}).$ (15)
This expression is strictly positive for all $x_{R}\in[0,1]$: it is linear in $x_{R}$ and takes the positive values $b_{R}>0$ at $x_{R}=0$ and $a_{F}>0$ at $x_{R}=1$. Hence $\dot{x}_{R}>0$ along this edge and $R$ eliminates $F$ in isolation.
The remaining edge calculations follow the same pattern. On the RRCC edge ($x_{F}=0$),
$\pi_{C}-\pi_{R}=a_{R}+x_{R}(b_{C}-a_{R})>0,$ (16)
so $C$ eliminates $R$. On the CCFF edge ($x_{R}=0$),
$\pi_{F}-\pi_{C}=a_{C}+x_{C}(b_{F}-a_{C})>0,$ (17)
so $F$ eliminates $C$. The induced boundary flow is therefore
$R\longrightarrow C\longrightarrow F\longrightarrow R,$ (18)
establishing a strict cyclic dominance structure and generating directed flow along the edges, with each edge carrying trajectories from one vertex to the next. This directed boundary cycle organizes the global dynamics and sets the stage for the heteroclinic cycle that follows.
To determine whether this boundary cycle governs the interior, we next analyze the interior equilibrium. This interior equilibrium $x^{\ast}$ is characterized by
$\pi_{R}(x^{\ast})=\pi_{C}(x^{\ast})=\pi_{F}(x^{\ast}),\qquad x_{R}^{\ast}+x_{C}^{\ast}+x_{F}^{\ast}=1.$ (19)
Because the payoff functions in (13) are linear in $x$, solving (19) yields a unique interior point. For strict cyclic-dominance matrices such as $A^{(1)}$, all coordinates of this solution are positive (Zeeman1980; Bomze1983; HofbauerSigmund1998).
For three-strategy replicator systems with a cyclic dominance structure, the interior equilibrium is a focus (Zeeman1980; Bomze1983; HofbauerSigmund1998). Consequently, the determinant of the Jacobian of the two-dimensional system on the simplex is positive, and stability is determined by the sign of the trace. The trace is
$\operatorname{tr}(J^{\ast})=\frac{a_{C}a_{F}a_{R}-b_{C}b_{F}b_{R}}{D},$ (20)
where $D>0$ is a positive polynomial in the payoff parameters that arises when solving (19) for $x^{\ast}$ and substituting into the Jacobian; its explicit form is not required here because only its sign matters, and positivity follows from the strict cyclic-dominance structure of $A^{(1)}$. This expression follows from standard linearization formulas for interior equilibria in three-strategy replicator systems (HofbauerSigmund1998). When
$a_{C}a_{F}a_{R}>b_{C}b_{F}b_{R},$ (21)
the interior equilibrium is repelling, and trajectories are pushed outward toward the directed boundary flow (18). The same inequality is equivalent to
$\rho_{R}\rho_{C}\rho_{F}=\frac{a_{F}}{b_{C}}\cdot\frac{a_{R}}{b_{F}}\cdot\frac{a_{C}}{b_{R}}>1,$ (22)
where $\rho_{R},\rho_{C},\rho_{F}$ are the eigenvalue ratios defined below, and together these conditions imply that trajectories approach a heteroclinic cycle connecting the three vertices (Zeeman1980; Bomze1983; HofbauerSigmund1998).
In this heteroclinic regime, the time the system spends near each vertex depends on the ratios of stable to unstable eigenvalues (Hofbauer1994). At $e_{F}=(0,0,1)$, the contraction rate along the $C$ direction is $a_{C}$ and the expansion rate along the $R$ direction is $b_{R}$, giving $\rho_{F}=a_{C}/b_{R}$. Similarly, $\rho_{R}=a_{F}/b_{C}$ and $\rho_{C}=a_{R}/b_{F}$. If
$\rho_{F}>\rho_{R}\quad\text{and}\quad\rho_{F}>\rho_{C},$ (23)
then we say the heteroclinic cycle is $F$-dominated: along the cycle, trajectories spend an asymptotically larger fraction of each successive excursion in neighborhoods of $e_{F}$ than near the other vertices, so sufficiently long empirical time averages can be made arbitrarily close to the fascist vertex, even though the trajectory itself does not converge to that vertex. This outcome is generated entirely by the payoff sign structure: although resistance defeats fascism in isolation, the presence of conflict-averse centrists produces a global cycle in which fascism dominates long stretches of the trajectory.
We now examine a second sign configuration in which resistance still defeats fascism in isolation, but centrism and fascism are mutually invadable along the CCFF edge. This situation arises when centrists gain a benefit from interacting with fascists, while fascists simultaneously exploit centrist norms. These behavioral assumptions are encoded in the diagonal-normalized payoff matrix
$A^{(2)}=\begin{pmatrix}0&-a_{R}&b_{R}\\ b_{C}&0&c_{C}\\ -a_{F}&c_{F}&0\end{pmatrix}$, (24)
where $a_{R},a_{F},b_{R},b_{C},c_{C},c_{F}>0$. Here $c_{C}>0$ captures the advantage centrists obtain from interacting with fascists, and $c_{F}>0$ captures the advantage fascists obtain from interacting with centrists. The signs $A_{RF}=b_{R}>0$ and $A_{FR}=-a_{F}<0$ preserve $R\succ F$, while $A_{CR}=b_{C}>0$ and $A_{RC}=-a_{R}<0$ preserve $C\succ R$. The key change is that the CCFF edge now exhibits mutual invadability rather than strict dominance.
To understand the dynamics, we first restrict attention to the CCFF edge ($x_{R}=0$). Along this edge,
$\pi_{C}=c_{C}x_{F},\qquad\pi_{F}=c_{F}x_{C},$ (25)
so each strategy invades the other’s vertex: $\pi_{C}>\pi_{F}$ near $x_{F}=1$ and $\pi_{F}>\pi_{C}$ near $x_{C}=1$. The resulting one-dimensional replicator system has a unique interior equilibrium. Setting $\pi_{C}=\pi_{F}$ gives $c_{C}x_{F}=c_{F}x_{C}$, and together with $x_{C}+x_{F}=1$ yields
$x_{R}^{\dagger}=0,\qquad x_{C}^{\dagger}=\frac{c_{C}}{c_{C}+c_{F}},\qquad x_{F}^{\dagger}=\frac{c_{F}}{c_{C}+c_{F}}.$ (26)
Standard two-strategy replicator dynamics imply that this point is locally stable and globally attracting on the CCFF edge (HofbauerSigmund1998; TaylorJonker1978).
Resistance can invade this edge equilibrium only if its per-capita growth rate when rare is positive. This requires its payoff against the mixture $(x_{C}^{\dagger},x_{F}^{\dagger})$ to exceed the average payoff in the resident population. Because $C$ and $F$ are at equilibrium on their shared edge, they have equal payoffs, so the resident average payoff is the common value
$\pi_{C}(x^{\dagger})=\pi_{F}(x^{\dagger})=c_{C}x_{F}^{\dagger}.$ (27)
Resistance’s payoff against this mixture is
$\pi_{R}(x^{\dagger})=(-a_{R})x_{C}^{\dagger}+b_{R}x_{F}^{\dagger}.$ (28)
Thus resistance can invade only if
$\pi_{R}(x^{\dagger})>c_{C}x_{F}^{\dagger}.$ (29)
Negating this yields the condition under which invasion fails:
$\pi_{R}(x^{\dagger})\leq c_{C}x_{F}^{\dagger},$ (30)
which states that the centrist–fascist coalition imposes a sufficiently strong combined effect on resistance.
Substituting the equilibrium coordinates
$x_{C}^{\dagger}=\frac{c_{C}}{c_{C}+c_{F}},\qquad x_{F}^{\dagger}=\frac{c_{F}}{c_{C}+c_{F}},$ (31)
into (30) and simplifying gives the weak inequality
$b_{R}\leq c_{C}\!\left(1+\frac{a_{R}}{c_{F}}\right).$ (32)
This is the precise threshold at which resistance’s growth rate becomes non-positive. For the coalition to be strictly uninvadable, we impose the strict version
$b_{R}<c_{C}\!\left(1+\frac{a_{R}}{c_{F}}\right).$ (33)
Here $b_{R}$ measures resistance’s advantage against fascists; $a_{R}$ captures the penalty centrists impose on resistance; and $c_{C}$ and $c_{F}$ represent the mutual benefits within the CCFF coalition. Notably, increasing $c_{C}$ enlarges the parameter region in which resistance cannot invade, making invasion harder, whereas increasing $c_{F}$ shrinks that region, making invasion easier.
In what follows, we focus on the parameter region in which
$c_{C}>b_{R}.$ (34)
Substantively, this corresponds to environments in which centrists’ perceived benefit from interacting with fascists exceeds resistance’s advantage in direct contests with fascists. This assumption automatically satisfies the anti-invasion condition, since
$c_{C}\!\left(1+\frac{a_{R}}{c_{F}}\right)>c_{C}>b_{R},$ (35)
so resistance’s growth rate at $x^{\dagger}$ is strictly negative.
Any interior rest point of the replicator dynamics would require all strategies to earn equal payoff; in particular, setting $\pi_{R}=\pi_{C}$ yields the necessary condition
$b_{C}x_{R}+a_{R}x_{C}+(c_{C}-b_{R})x_{F}=0.$ (36)
When $c_{C}>b_{R}$, all coefficients in this expression are strictly positive for every $x_{i}>0$, so the equality cannot hold in the interior of the simplex and no interior equilibrium exists. On the remaining boundary, the RRCC and RRFF edges each support monotone flow (with $C$ eliminating $R$ on the RRCC edge and $R$ eliminating $F$ on the RRFF edge), and all three vertices are unstable. With $x^{\dagger}$ the only non-vertex equilibrium and no attracting periodic orbits in the planar replicator dynamics on the simplex, classical convergence results (HofbauerSigmund1998) imply that every interior trajectory converges to $x^{\dagger}$. In this second configuration, the behavioral asymmetry between centrists and fascists produces a stable coalition that excludes resistance and maintains a persistent, positive equilibrium frequency of fascism.
Together, these results show how empirically grounded payoff sign patterns generate distinct global behaviors. In the first configuration, conflict aversion creates a dominance cycle that drives the system toward a heteroclinic orbit on which the flow can spend an asymptotically disproportionate amount of time near the fascist vertex. In the second configuration, mutual invadability along the CCFF edge produces a stable coalition that excludes resistance, allowing fascism to persist at potentially high frequency.
5 Discussion and Conclusions
This minimal evolutionary model captures a politically salient mechanism: actors who frame social conflict primarily as a problem of polarization can unintentionally undermine resistance and thereby facilitate the persistence of authoritarian projects. The model formalizes how conflict aversion, civility norms, and reputational dynamics—well documented in political science and political psychology—can reshape strategic incentives in morally asymmetric conflicts. Across the two empirically grounded mechanisms generated by our payoff sign patterns, introducing a conflict-averse strategy alters outcomes in ways that neither two-strategy reasoning nor symmetric models of polarization can capture.
The first mechanism shows that even when resistance is intrinsically stronger than authoritarianism in direct confrontation, the presence of conflict-averse actors can generate a heteroclinic cycle. In parameter regions where the eigenvalue ratios favor fascism, this cycle produces a regime in which fascism repeatedly resurges and can dominate long stretches of the trajectory. In those same regions—where trajectories spend disproportionately long near the fascist vertex—this structure offers a formal account of how civility norms, media framing of protest as disruptive, and public discomfort with confrontation can repeatedly regenerate authoritarian strength. The resulting cycling implies that democratic movements may experience recurring episodes in which authoritarian actors come close to dominating the population, even though resistance would eliminate authoritarianism were centrists absent. Conflict aversion thus does more than impede resistance: it reorganizes the global dynamics in ways that prevent stable democratic dominance.
The second mechanism reveals a distinct pathway through which conflict aversion can advantage authoritarian projects. When centrists impose penalties on resistance that are sufficiently strong—relative to resistance’s advantage against fascism and to the coalition parameters $c_{C}$ and $c_{F}$ that govern mutual benefits within the centrist–fascist coalition, as captured by the invasion condition $b_{R}<c_{C}(1+a_{R}/c_{F})$—the two may form a dynamically stable coalition along the CCFF edge that resistance cannot invade. Notably, increasing $c_{C}$ raises the right-hand side of this inequality, expanding the parameter region in which invasion fails, whereas increasing $c_{F}$ lowers it, shrinking that region. Although resistance still defeats fascism in isolation, the alliance between conflict-averse actors and authoritarian actors can lock the system into a boundary equilibrium with a persistent and potentially high fascist share. The coalition’s ability to exclude resistance depends jointly on the cost $a_{R}$ that centrists impose on resistance, resistance’s advantage $b_{R}$ against fascists, and the mutual reinforcement parameters $c_{C}$ and $c_{F}$, while the equilibrium’s location along the CCFF edge—and thus how close it lies to the fascist vertex—is determined solely by $c_{C}$ and $c_{F}$. This mechanism formalizes how appeals to civility, moderation, or both-sides norms can unintentionally sustain authoritarian projects by generating a stable alignment that structurally disadvantages democratic resistance.
Taken together, the two mechanisms show that the strategic role of centrism is not neutral. Norms and behaviors that discourage confrontation may erode democratic defenses by preventing resistance from achieving stable dominance. From a normative perspective, framing asymmetric conflicts as problems of polarization—treating norm-violating and norm-defending behavior as symmetrically problematic—can inadvertently advantage authoritarian actors. The model clarifies that focusing solely on pairwise matchups (for example, whether resistance defeats authoritarianism in isolation) obscures how third-party reactions reshape strategic landscapes.
The model is intentionally stylized. It abstracts away from institutions, media ecosystems, social networks, and spatial heterogeneity, focusing instead on the core strategic logic of interactions among three behavioral types. Strategies are fixed rather than learned or adapted, and payoffs are deterministic rather than stochastic. Nonetheless, even this bare-bones representation shows that treating centrism as a distinct behavioral strategy—rather than as an ideological midpoint or omitting it entirely—can yield conclusions about authoritarian persistence that differ fundamentally from those obtained in two-strategy or symmetric polarization models. The central insight is that in morally asymmetric conflicts, behaviors that prioritize conflict avoidance can inadvertently reshape strategic incentives in ways that hinder democratic resistance and enable authoritarian persistence.
Acknowledgments
The author used Chat-GPT 5.1 to refine and tighten text and to proofread this manuscript.